OpenStudy (anonymous):

A rectangle has length 12 feet and width 8 feet. Every dimension of the rectangle is multiplied by 3 over 4 to form a similar rectangle. How is the ratio of the areas related to the ratio of corresponding sides?

OpenStudy (kkutie7):

3 over 4 like this \[\frac{3}{4}\] well you can just look at the equation for the area of a rectangle. \[A=L*W\] it tells you that the dimensions are being multiplied by the fraction so \[A=\frac{3}{4}L*\frac{3}{4}W\] you can factor out the fraction \[A=\frac{3}{4}(L*W)\] so when the dimensions are multiplied by something so is the area \[\frac{3}{4}(L*W)=\frac{3}{4}(A)\]

OpenStudy (kkutie7):

The last thing there isn't "solved" its just a way to think about it

OpenStudy (anonymous):

Thank you!! :)